OpenStudy (anonymous):

Find all possible solutions for x such that ∆ABC is congruent to ∆DEF. ∆ABC: sides length 7, 9, x ∆DEF: sides length 7, 10, x-1 The solution to x = _______.

OpenStudy (anonymous):

Since the triangles are congruent by sides, the sides are equal. We know that in triangle ABC, the sides are 7, 9, and x. In triangle DEF the sides are 7,10, and x-1.

OpenStudy (anonymous):

Find all possible solutions for x and y such that ∆ABC is congruent to ∆DEF. ∆ABC: sides length 18, 25, and 6x – 4 ∆DEF: sides length 25, 38, and 5 – y ∆ABC ≅ ∆DEF when x is _________ and y is -------- .

OpenStudy (anonymous):

Since it is congruent, the sides must be equal to each other, we know that 7=7, but 9 can't equal 10, so 9=x-1. if we solve for x, x=10

OpenStudy (anonymous):

we apply the same thing here, in the two triangles all of the sides must be equal. In both triangles they have the same side length of 25, since 18 and 38 are not equal to each other, they must be equal to the other corresponding sides. 18=5-y and 38=6x-4

OpenStudy (anonymous):

solve for x and y

OpenStudy (anonymous):

understand? or nah?

OpenStudy (anonymous):

add 5 divide by Y add 4 divide by 6 ????

OpenStudy (anonymous):

subtract 5

OpenStudy (anonymous):

you would subtract 5 and divide by -1. because you want y to become posititve. and yes you would add 4 and divide by 6.

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

Very much!

OpenStudy (anonymous):

Your welcome then (;

OpenStudy (anonymous):

Thanks read your testimonial

OpenStudy (anonymous):

I dont know what a testimonial is lol